# An evaluation of CFEAR Radar Odometry

### v1

@@ Abstract
This article describes the method CFEAR Radar odometry, submitted to a competition at the Radar in Robotics workshop, ICRA 2024111https://sites.google.com/view/radar-robotics/competition. CFEAR [1] is an efficient and accurate method for spinning 2D radar odometry that generalizes well across environments. This article presents an overview of the odometry pipeline with new experiments on the public Boreas dataset. We show that a real-time capable configuration of CFEAR – with its original parameter set in [2] – yields surprisingly low drift in the Boreas dataset. Additionally, we discuss an improved implementation and solving strategy that enables the most accurate configuration to run in real-time with improved robustness, reaching as low as 0.66% translation drift at a frame rate of 68 Hz. A recent release of the source code is available to the community https://github.com/dan11003/CFEAR_Radarodometry_code_public, and we publish the evaluation from this article on https://github.com/dan11003/cfear_2024_workshop.
@@ II CFEAR Radarodometry
Motion is incrementally estimated for the latest sweep by finding the pose $\mathbf{x}^{t}\in SE(2)$ that minimizes:
@@ V Evaluation
Surprisingly CFEAR-CTF-S10 reached as low as 0.66% in the Boreas training set. In this experiment, the coarse-to-fine strategy yields only a minor improvement which we attribute to the otherwise rare failures with the used parameters. Larger improvements were observed when experimenting with other parameter sets. The major improvement is attributed to the use of additional keyframes which is possible with the significantly more efficient nearest neighbor search.
@@
Table I: Evaluation for training and test sequences in the Boreas dataset, and 8 sequences from the Oxford training set. Drift is measured by the translation error [%], rotation error (deg/100m). In Boreas, CFEAR-3 runs at 130.5 Hz, CFEAR-CTF at 119.8 Hz, and CFEAR-CTF-S10 at 68.0 Hz. Results from test sequences will be reported when available.
@@
Sequence(s) (transl. error [%] / rot. error [deg/100m])
Training set (0.83/0.35) (0.79/0.35) (0.66/0.34)
2020-12-04-14 - - -
2021-01-26-10 - - -
2021-02-09-12 - - -
2021-03-09-14 - - -
2021-04-22-15 - - -
2021-06-29-18-53 - - -
2021-06-29-20-43 - - -
2021-09-08-21 - - -
2021-09-09-15 - - -
2021-10-05-15 - - -
2021-10-26-12 - - -
2021-11-06-18 - - -
2021-11-28-09 - - -
2021-11-28-09 - - -
Test set - -
@@ VI Conclusion
This work presented CFEAR, and discussed both robustness and efficiency improvements. An evaluation was carried out on the Boreas dataset, reaching as low as 0.66 % drift in the training set. Additional experiments in the Oxford, and MulRan dataset demonstrate a high level of generalization, with 1.16% and 1.18% drift respectively without tuning any parameters. Future work ought to investigate the cause of systematic trajectory errors. Results from test sequences will be imported when available.

### v2

@@ Abstract
This article describes the method CFEAR Radar odometry, submitted to a competition at the Radar in Robotics workshop, ICRA 2024111https://sites.google.com/view/radar-robotics/competition. CFEAR [1] is an efficient and accurate method for spinning 2D radar odometry that generalizes well across environments. This article presents an overview of the odometry pipeline with new experiments on the public Boreas dataset. We show that a real-time capable configuration of CFEAR – with its original parameter set in [2] – yields surprisingly low drift in the Boreas dataset. Additionally, we discuss an improved implementation and solving strategy that enables the most accurate configuration to run in real-time with improved robustness, reaching as low as 0.61% translation drift at a frame rate of 68 Hz. A recent release of the source code is available to the community https://github.com/dan11003/CFEAR_Radarodometry_code_public, and we publish the evaluation from this article on https://github.com/dan11003/cfear_2024_workshop.
@@ I Introduction
Recent work on radar odometry [4, 5] has investigated the benefit of fusing IMU or Doppler information, which can be particularly useful in geometrically degenerate cases. In contrast, the odometry pipeline presented here, namely CFEAR, operates on range-intensity measurements without additional sensors. An estimated trajectory with the method presented in this submission is presented in Fig. 1.
@@ II CFEAR Radarodometry
Motion is incrementally estimated by finding the pose $\mathbf{x}^{t}\in SE(2)$ that minimizes:
@@ V Evaluation
Surprisingly CFEAR-CTF-S10 reached as low as 0.66% in the Boreas training set, and 0.61% in the test set with ground truth held out. In this experiment, the coarse-to-fine strategy yields only a minor improvement (roughly 5%) which we attribute to the otherwise rare failures with the used parameters. Larger improvements were observed when experimenting with other parameter sets. The major improvement is attributed to additional keyframes which is feasible with the more efficient nearest neighbor search.
@@
Table I: Evaluation for training and test sequences in the Boreas dataset, and 8 sequences from the Oxford training set. Drift is measured by the translation error [%], rotation error (deg/100m). In Boreas, CFEAR-3 runs at 130.5 Hz, CFEAR-CTF at 119.8 Hz, and CFEAR-CTF-S10 at 68.0 Hz.
@@
Dataset (transl. error [%] / rot. error [deg/100m])
Boreas training set (0.83/0.35) (0.79/0.35) (0.66/0.34)
Boreas test set - - (0.61/0.2)
@@ VI Conclusion
This work presented CFEAR, and discussed both robustness and efficiency improvements. An evaluation was carried out on the Boreas dataset, reaching as low as 0.61 % drift in the test set with ground truth held out. Additional experiments in the Oxford, and MulRan dataset demonstrate a high level of generalization, with 1.16% and 1.18% drift respectively – without tuning any parameters. Future work ought to investigate the cause of systematic trajectory errors.